M1_IND2. An online self-publishing company estimates that the fixed costs of its first major project will be $41,800, the variable cost will be $18 per book, and the selling price per book will be $23.50. a) How many books must be sold to break even? b) What is the total revenue at the break-even point? c) If the editors take a total of $35,200 in salary (what cost is affected?) – how many books must be sold to break even? d) What is the total revenue for the break-even volume when the publishers take a total of $35,200 in salary?
Answer #1

a)break even point =fixed cost/marginal profit=fixed cost/(selling cost-variable cost)

=41800/(23.5-18)=7600 books

b)

total revenue =7600*23.5=178600

c)

break even point=(41800+35200)/(23.5-18)=14000

d)

total revenue=14000*23.5=329000